Assume that banks do not hold excess reserves. Banking system has $50 million in reserves and a reserve requirement of 10%. Public holds 20 million in currency . Then the public decides to withdraw $5 million in currency from the banking system. If the banking system wants to keep money supply stable by changing the reserve requirement. What will the new reserve requirement be?

A)8.1%

B)9.1%

C)9.7%

D) 10%
Answer #1

Banking system has $50 million in reserves and a reserve requirement of 10%. This implies that the system has a total of 50/0.1 = 500 million deposits. Public holds 20 million in currency . This makes the money supply equal to 500 + 20 = 520 million and monetary base = 50 + 20 = 70 million. Money multiplier is Money supply / monetary base = 520/70 = 7.428571

Then the public decides to withdraw $5 million in currency from the banking system. This makes the deposits equal to 495 billion and currency equal to 25 million. To keep money supply unchanged at 520 million, multiplier should be same at 7.428571 and so monetary base should be 70 million. Now currency is 25 million so reserves should be 70 - 25 = 45 million. Hence new reserve requirement is 45 / 495 = 9.1%

Select 9.1%.